FTIR OBSERVATION VIBRATIONAL FUNDAMENTALS OF SiCH AND H_{2}SiC_{2}H IN SOLID Ar

DONGSHENG HAN, C.M.L. RITTBY AND \underline{W.R.M. GRAHAM}, Department of Physics, Texas Christian University, Fort Worth, TX 76129.

Fourier transform infrared studies have resulted in the first observation of vibrational spectra of the linear SiCH and non-planar H_{2}SiCCH radicals produced by trapping the products of the vacuum ultraviolet photolysis of a mixture of SiH_{4} silane, and C_{2}H_{2} acetylene, in an Ar matrix at about 10 K. The silicon-carbon stretching mode ν_{1}(σ) of SiCH has been observed at 1010.4 cm^{−1}. Two vibrational fundamentals of H_{2}SiC_{2}H have been identified, the ν_{3}(a’), siliconcarbon stretching mode at 2055.6 cm^{−1} and the ν_{4}(a’), hydrogen-silicon bending mode at 926.8 cm^{−1}. The results of extensive measurements of D and ^{13}C isotopic shifts are in good agreement with the predictions of coupled cluster ab initio calculations performed at the CCSD(T)/6-31G** level and confirm the proposed identifications.

aResearch supported by the Welch Foundation